(A) How many different ways are there for 17 of the 24 people to develop the disease? (B) Now suppose 9 of the 24 people carry a specific genetic mutation, and 5 of those are in the group that develops the disease. How many different ways are there for this to happen?
Answer #1

A)

C(24,17)

C(n,r)=C(24,17)C(n,r)=C(24,17)

= 346104

B)

C(24,9)*C(9,5) = 1307504*126 = 164745504